### v3

# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

Abstract
We introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter within the critical strip. We establish a rigidity phenomenon: Two solutions with the same initial condition equal to 11, corresponding to two different parameters lying on a horizontal line in the critical strip, cannot both be bounded on $1,+\infty)$. In particular, when $\Re(w)\neq\tfrac{1}{2}$, the two solutions corresponding respectively to the parameters $w$ and $1-\bar{w}$, both taken with initial condition 11, cannot be simultaneously bounded. This result provides an answer to the Dynamical Conjecture formulated in [2].
Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
MSC subject classifications: 34A30, 34E05
1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equation
$\dot{x}=w\,t^{-1}x+(1-w)\,t^{-1}\eta(t),$ (1)
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
where the parameter $w\in\mathbb{C}$. The the nonhomogeneous term $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function satisfying the following hypothesis (H):
$|\eta(t)|<1,\quad\forall t\geq 1,\quad\text{and}\quad\exists\rho\in\mathbb{C}^{*}:\ \sup_{t\geq 1}\left|\int_{1}^{t}(\eta(u)-\rho)\,du\right|<+\infty.$ (H)
We call $\rho$ the rotation number of the function $\eta$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. More precisely, our aim is to prove that for every $w,w^{\prime}\in\mathbb{C}$ with $\Re(w)\neq\Re(w)$ and $\Im(w)=\Im(w^{\prime})$, the solution $x_{w}$ and the solution $x_{w^{\prime}}$ with initial conditions $x_{w}(1)=x_{1-\overline{w}}(1)=1$ cannot both be bounded on $1,+\infty)$. In the study that follows, we fix a specific function $\eta$ which is locally integrable satisfying hypothesis (H) with a fixed rotation number $\rho\in\mathbb{C}^{*}$. The function
$t\mapsto t^{w}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
is continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
$\psi_{w}(z,t)=t^{w}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
The analysis in this manuscript is valid only when the real part of the parameter is in $(0,1)$. We consider the following parameter set,
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)\in(0,1),\quad\Im(w)\neq 0\Big\}.$
Let us introduce some notations.
Notation 1.
For every $w\in B$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
Notation 2.
We denote by $\mu$ the function $\mu:B\mapsto\mathbb{C}$ defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $w\in B$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
$\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
where the function $\mu$ is given by the Notation 2. Further,
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
Proof.
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions are oscillating and diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $z=\mu(w)$. To prove that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$, observe that by definition of $\mu(w)$,
$\mu(w)+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$
Using equation (2), we then obtain
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{w}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
From the Hypothesis (H) we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
Now, prove the last item of the present Lemma. From equation (3), we have
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w)t^{{w}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,\quad\forall t\geq 1.$
Using the integration by parts formula, we obtain
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w^{2})t^{w}\int_{t}^{+\infty}u^{-2-w}\int_{1}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\right|\leq\,t^{-1}\dfrac{|1-w^{2}|}{1+\Re(w)}\sup_{t\geq u}\Big|\int_{1}^{u}(\eta(v)-\rho)dv\Big|.$
By the Hypothesis (H), we obtain
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
The following theorem, show in particular, that if $\Re(s)\neq\frac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Consider the function $\mu$ given by the Notation 2. For every $s,s^{\prime}\in B$ with $\Re(s)\neq\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$, we have
$\big(\mu(s),\,\mu(s^{\prime}))\neq(1,1).$
In particular, for every $s\in B$ we have the following implication:
$\big(\mu(s),\,\mu(1-\overline{s})\big)=(1,1)\implies\Re(s)=\frac{1}{2}.$
2 Main Proposition
From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We show this in the following proposition.
Proposition 5.
Let be $s:=\sigma+i\tau\in B$ and $s^{\prime}:=\sigma^{\prime}+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. Suppose that $\sigma>\sigma^{\prime}$, then
$\Re\big(t^{-i\,\tau}\psi_{s}(1,t)\big)>\Re\big(t^{-i\,\tau}\psi_{s^{\prime}}(1,t)\big),\qquad\forall t>1.$
Proof.
Let be $w\in B$ and denote $w:=\alpha+i\beta$ . For every $t\geq 1$ define
$\psi_{w}(t):=t^{-i\beta}\psi_{w}(1,t).$
According to Notation 1, these functions satisfy the differential equations
$\dot{\psi}_{w}(t)=\alpha\,t^{-1}\psi_{w}(t)+(1-w)\,t^{-1-i\beta}\eta(t),\quad\psi_{w}(1)=1.$
In order to simplify the notation, denote
$x_{w}(t):=\Re\big(\psi_{w}(t)\big)-1,\qquad\forall t\geq 1.$
The previous differential equations, implies that
$\dot{x}_{w}(t)=\alpha\,t^{-1}x_{w}(t)+\alpha t^{-1}+t^{-1}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=0.$ (4)
Let be $s:=\sigma+i\tau\in B$ and $s:=\sigma^{\prime}+i\tau\in B$ such that $\sigma>\sigma^{\prime}$. The previous differential equation implies
$\dot{x}_{s}(t)-\dot{x}_{s^{\prime}}(t)$ $=\sigma\,t^{-1}x_{s}(t)-\sigma^{\prime}\,t^{-1}x_{s^{\prime}}(t)$ (5)
$+(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\,t^{-i\tau}\,\eta(t)\big)\big).$
We have $x_{s}(1)=x_{s^{\prime}}(1)=0$. The previous differential equation implies that
$\dot{x}_{s}(1)-\dot{x}_{s^{\prime}}(1)=(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\eta(t)\big)\big).$
By the hypothesis (H), we have $|\eta(t)|<1$ for all $t\geq 1$. By hypothesis we have $\sigma>\sigma^{\prime}$, then $\dot{x}_{s}(1)>\dot{x}_{s^{\prime}}(1)$. By continuity there exists $\epsilon>0$ such that
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,1+\epsilon).$
Define the time $t_{*}>1$ as
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ x_{s}(u)>x_{s^{\prime}}(u)\}.$ (6)
The strategy to prove the present Proposition is to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$x_{s}(t_{*})=x_{s^{\prime}}(t_{*}),$
and we have
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,t_{*}).$ (7)
We have two cases: $x_{s}(t_{*})\geq 0$ and $x_{s}(t_{*})<0$.
If $x_{s}(t_{*})\geq 0$, then $x_{s^{\prime}}(t_{*})\geq 0$. Using the fact that $\sigma>\sigma^{\prime}$, from the differential equations (5), we get
$\dot{x}_{s}(t_{*})-\dot{x}_{s^{\prime}}(t_{*})$ $=(\sigma-\sigma^{\prime})t_{*}^{-1}x_{s}(t_{*})$
$+(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big)$
$\geq(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
By the hypothesis (H), we obtain $\dot{x}_{s}(t_{*})>\dot{x}_{s^{\prime}}(t_{*})$. There exists $t<t_{*}$ close enough to $t_{*}$ such that $x_{s}(t)<x_{s^{\prime}}(t)$. We have obtained a contradiction with equation (7).
If $x_{s}(t_{*})<0$, then $x_{s^{\prime}}(t_{*})<0$. By continuity there exists $\tilde{\epsilon}>0$ such that
$0>x_{s}(u)>x_{s^{\prime}}(u),\quad\forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
Implies
$\delta_{s,s^{\prime}}(u):=x_{s}(u)x_{s^{\prime}}(u)-(x_{s}(u))^{2}=x_{s}(u)(x_{s^{\prime}}(u)-x_{s}(u))>0,\ \forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
and $\delta_{s,s^{\prime}}(t_{*})=0$. Then we must find $\dot{\delta}_{s,s^{\prime}}(t_{*})\leq 0$. As the previous technique, prove that in our system that $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. From the differential equation (4), we have
[…]
$x_{s^{\prime}}(t)\dot{x}_{s}(t)$ $=\sigma\,t^{-1}x_{s^{\prime}}(t)x_{s}(t)+\sigma t^{-1}x_{s^{\prime}}(t)+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big),$
$x_{s}(t)\dot{x}_{s^{\prime}}(t)$ $=\sigma^{\prime}\,t^{-1}x_{s}(t)x_{s^{\prime}}(t)+\sigma^{\prime}t^{-1}x_{s}(t)+t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big).$
Implies
$\dot{\delta}_{s,s^{\prime}}(t)$ $=2\sigma\,t^{-1}{\delta}_{s,s^{\prime}}(t)+t^{-1}\big(\sigma x_{s^{\prime}}(t)+\sigma^{\prime}x_{s}(t)\big)$
$+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big)$
$-t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big)$
$-2\sigma t^{-1}x_{s}(t).$
At the point $t_{*}$ we get
$\dot{\delta}_{s,s^{\prime}}(t_{*})$ $=-t_{*}^{-1}\big(\sigma-\sigma^{\prime}\big)x_{s}(t_{*})\big(1-\Re\big(t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
Use the fact that $\sigma>\sigma^{\prime}$ and $x_{s}(t_{*})<0$, we obtain $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. Same, we obtained a contradiction. Then $t_{*}=+\infty$. ∎
3 Proof of the Theorem 4
Intuitively, the solution $\psi_{w}(1,\cdot)$cannot be bounded when $\Re(w)\in\left(\tfrac{1}{2},1\right)$. Indeed, if $\Re(w)\in(0,1)$, then |w|>|1w||w|>|1-w|. Since the initial condition satisfies $\psi_{w}(1,1)=1$ and $\sup_{t\geq 1}|\eta(t)|\leq 1$, for $t$ close to 11 the homogeneous part dominates the non-homogeneous part in differential equation (1). This may therefore lead to the unboundedness of the solution $\psi_{w}(1,\cdot)$. In this manuscript, we compare and study the boundedness of two solutions corresponding to two complex parameters with the same imaginary part, in order to eliminate the imaginary part (which may be very large) from the computations.
Proof of the Theorem 4.
We prove the theorem by contradiction. Suppose there exist $s,s^{\prime}\in B$ with $\Re(s)>\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$ such that $\big(\mu(s),\,\mu(s^{\prime})\big)=(1,1)$. The solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) are both bounded on $1,+\infty)$. Thanks to the second Item of the Lemma 3, we have
$\sup_{t\geq 1}\Big|t\Big(\psi_{s}(1,t)+\rho\frac{1-s}{s}\Big)\Big|<+\infty\ \text{and}\ \sup_{t\geq 1}\Big|t\Big(\psi_{s^{\prime}}(1,t)+\rho\frac{1-s^{\prime}}{s^{\prime}}\Big)\Big|<+\infty.$
Use the Proposition 5, we get
$\lim_{t\to+\infty}\Re\Big(t^{-i\Im(s)}\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)\Big)\geq 0.$
Since $s,s^{\prime}\in B$, then $\Im(s)\neq 0$, that gives
$\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)=0.$
By the hypothesis (H), we have $\rho\neq 0$. then
$\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}=0$
We have obtained a contradiction with $\Re(s)\neq\Re(s^{\prime})$. ∎
4 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$ (the fractional part of $t$) is a locally integrable function and satisfies $0\leq\eta(t)<1$ with a rotation number $\rho$ (here $\rho=1/2$). The Hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
In other words, by considering the integral representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have the relations
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (8)
Thanks to the Riemann functional equation [1, Theorem 2.1, p. 13], when $\zeta(s)=0$ then $\zeta(1-\bar{s})=0$. Implies $\big(\mu(s),\ \mu(1-\bar{s})\big)=(1,1)$. By the Theorem 4 we deduce that $\Re(s)=\Re(1-\bar{s})$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.