### v4

# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

Abstract
In this manuscript, we introduce a family of parametrized nonhomogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a suitable rigidity hypothesis under which a phenomenon of structural asymmetry is established: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
MSC subject classifications: 34A30, 34E05
1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equations
$\dot{x}=\Re(w)\,t^{-1}x+(1-w)\,t^{-1-i\,\Im(w)}\eta(t),$ (1)
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
where $w\in\mathbb{C}$ and $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. This result provides an answer to the Dynamical Conjecture formulated in [2]. In the study that follows, we fix a specific function $\eta$ which is locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
is continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
$\psi_{w}(z,t)=t^{\Re(w)}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
Let us introduce some notations.
Notation 1.
For every $w\in\mathbb{C}$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1) given by equation (2). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
Notation 2.
Denote by $B\subset\mathbb{C}$ the following set
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
Let $\mu$ denote the function $\mu:B\mapsto\mathbb{C}$, defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $w\in B$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
$\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
where the function $\mu$ is given by the Notation 2.
Proof.
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $z=\mu(w)$. To prove that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$, observe that by definition of $\mu(w)$,
$\mu(w)+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$
Using equation (2), we then obtain
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
By Hypothesis we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We show this in the following proposition.
Proposition 4.
Let be $s:\sigma+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. The function $\delta_{s}$ defined as
$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
satisfies the following two equalities:
$\dot{\delta}_{s}(1)$ $=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big),$
${\delta}_{s}(t)$ $=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$\ +4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right),\qquad\forall t\geq 1.$
According to the notation 1 consider the following rigidity hypothesis H:
  • H.
    For every $\beta\in\mathbb{R}$ and every $t>1$,
    $\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right)>0.$
The previous hypothesis suggest the rigidity of the sign. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
The following theorem shows that, under the rigidity hypothesis H, if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 5.
Consider the function $\mu$ defined in Notation 2. Under hypothesis H, for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
$\big(\mu(s),\,\mu(1-\overline{s})\big)\neq(1,1).$
2 Proof of the Proposition 4 and Theorem 5
Proof of the Proposition 4.
Let be $w\in B$ and denote $w:=\alpha+i\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\Re\big(\psi_{w}(1,t)\big).$ (4)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+t^{-\frac{3}{2}}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=1.$ (5)
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(\frac{1}{2},1)$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\bar{s}}(t)\big)+t^{-\frac{3}{2}}\Re\Big((1-2i\tau)t^{-i\tau}\,\eta(t)\Big).$
Use the fact that $x_{s}(1)=x_{1-\bar{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2+\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
From the differential equation (5), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\bar{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2(\sigma-\frac{1}{2})t^{-1}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
$+(\sigma-\frac{1}{2})t^{-1}\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
Define
$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
By the notation (4), we have ${x_{s}}(t)-{x}_{1-\bar{s}}(t):=t^{-\frac{1}{2}}\delta_{s}(t)$. We obtained
$\dot{\delta}_{s}(t)=$ $\frac{1}{2}t^{-1}\delta_{s}(t)+(\sigma-\frac{1}{2})^{2}t^{-\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\delta_{s}(u)\,du$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
Since ${\delta}_{s}(1)=0$ and since by hypothesis we have $|\eta(1)|<1$, then
$\dot{\delta}_{s}(1)=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big).$
Now, prove the second item of the present Proposition. Integrate the previous equation, we obtain
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\,du\Big).$
Use the integration by pars formula to the last integral, we get
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big)$
$-4(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big).$
Equivalent to
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\big[(\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big]$
$-4(\sigma-\frac{1}{2})\Big[1+\Re\Big((1-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big)\Big].$
According to Notation 1, we obtain
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right).$
Proof of the Theorem 5.
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
$\delta_{s}(t):=\Re\big(\psi_{s}(1,t)-\psi_{1-\bar{s}}(1,t)\big).$
Thank to the Proposition 4, we have $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
$\delta_{s}(t)>0,\quad\forall t\in(1,1+\epsilon).$
Define the time $t_{*}>1$ as
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ \delta_{s}(u)>0\}.$
To prove of the present Proposition it is equivalent to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$\delta_{s}(t_{*})=0,$ (6)
and we have
$\delta_{s}(u)>0,\quad\forall u\in(1,t_{*}).$ (7)
Thank to the Proposition 4, we have
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right).$
Since $2\sigma-1>0$, thanks to the hypothesis H, we obtain
${\delta}_{s}(t)>\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$ (8)
We obtain
${\delta}_{s}(t_{*})>\frac{1}{2}\int_{1}^{t_{*}}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t_{*}}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,$
Since $2\sigma-1>0$, by equation (7) we get
${\delta}_{s}(t_{*})>0$
We obtained a contradiction with equation (6). Then $t_{*}=+\infty$. In other words, $\delta_{s}(t)>0$ for all $t>1$. The equation (8) implies
$\dot{\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{2}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$
That gives
${\delta}_{s}(t)>2(\sigma-\frac{1}{2})^{2}(t^{\frac{1}{2}}-1)\int_{1}^{2}u^{-\frac{3}{2}}{\delta}_{s}(u)du,\quad\forall t>1.$
Then $\lim_{t\to+\infty}\delta_{s}(t)=+\infty$. This implies that the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ are not both bounded on $1,+\infty)$. By Lemma 3, we deduce that
$(1,1)\neq\big(\mu(s),\,\mu(1-\overline{s})\big).$
More precisely, the proof of Proposition 4 shows that under the rigidity hypothesis H, for all $s\in B$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following equation
$\Re\big(\mu(s)\big)-\Re\big(\mu(1-\overline{s})\big)\leq 2\Re\big(\eta(1)\big).$
In this case, $\dot{\delta}_{s}(1)\leq 0$; That is, the function does not necessary start positively from 11. Hence the proof of the Theorem 5 cannot be continued.
3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will now see that hypothesis H, is also satisfied. From the equation (2), we have
$\psi_{w}(1,t)=t^{\Re(w)}\Big(1+(1-w)\int_{1}^{t}u^{-1-w}\,\{u\}\,du\Big),\quad\forall w\in\mathbb{C},\quad\forall t>1,$
which can be written as
$\psi_{w}(1,t)$ $=t^{\Re(c)}\Big[1+J(t,\lfloor t\rfloor)+\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)\Big],$
$J_{w}(x,n)$ $:=(1-w)\int_{n}^{x}u^{-1-w}\,(u-n)\,du,$
with $\sum_{n=1}^{\lfloor t\rfloor-1}J_{w}(n+1,n)=0$ when $t\in[1,2)$. Use the integration by parts formula, we obtain
$J_{w}(x,n)$ $=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,\frac{d}{du}\Big(u^{-1}\,(u-n)\Big)du$
$=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,u^{-2}du.$
For every $\beta\in\mathbb{R}$ and $t\geq 2$ we have
$\Big|t^{\frac{1}{2}}J_{\frac{1}{2}+i\,\beta}(x,n)-J_{i\,\beta}(x,n)\Big|$ $\leq\int_{n}^{x}\Big(\big(t^{\frac{1}{2}}x^{\frac{1}{2}}-x\big)+(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)\,\Big)u^{-2}du$
Implies
$\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)$ $\geq t^{\frac{1}{2}}-1-\int_{\lfloor t\rfloor-1}^{t}(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)u^{-2}du$
$+\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big(n+1+u)\,u^{-2}du$
$-t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big((n+1)^{\frac{1}{2}}+u^{\frac{1}{2}}\big)u^{-2}du$
$\geq\frac{1}{3}\big(t^{\frac{1}{2}}-1\big)+\left(1-\frac{1}{t}\right),\qquad\forall t\geq 1.$
We recall that $\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)=0$ for all $t\in[1,2)$. We deduce that
$\Re\Big(\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)\Big)>0,\quad\forall t>1.$
Then the Hypothesis hypothesis H is satisfied. By definition of the function $\mu$ in the Definition 2, we have
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
In other words, by considering the integral representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have the relations
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
Since the Hypothesis hypothesis H is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.