### v6

# Rigidity Hypothesis for the Structural Asymmetry of the Bounded Solutions of a Parametric Complex Linear Differential Equation

Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$. This rigidity hypothesis is satisfied when the non-homogeneous term is the fractional part function.
Keywords: Non-homogeneous complex linear differential equation, bounded solutions.
MSC subject classifications: 34A30, 34E05
1 Main results
In this manuscript, we study the following non-homogeneous linear complex differential equations:
$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta:[1,+\infty)\to\mathbb{C},$
where $w\in\mathbb{C}$ is the parameter and the non-homogeneous term $\eta$ is a normalized, locally integrable function satisfying $|\eta(t)|<1$ for all $t\geq 1$. We are interested in the initial conditions of the bounded solutions. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
$\dot{x}=\Re(w)t^{-1}x+(1-w)t^{-1-i\Im(w)}\eta(t),$ (1)
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C}.$
In the following study, we fix a specific function $\eta$ as defined above. The function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
is absolutely continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
$\psi_{w}(z,t)=t^{\Re(w)}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
Let us introduce some notations.
Notation 1.
For every $w\in\mathbb{C}$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1) given by equation (2). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
Notation 2.
Denote by $\mathbb{C}_{+}$ the right half-plane defined as
$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
Let $\mu$ denote the function $\mu:\mathbb{C}_{+}\mapsto\mathbb{C}$, defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in\mathbb{C}_{+}.$
The function $\mu$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
Lemma 3.
Let $w\in\mathbb{C}_{+}$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
$\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty\quad\text{if and only if}\quad z=\mu(w),$
where the function $\mu$ is given by the Notation 2.
Proof.
Let $w\in\mathbb{C}_{+}$ be fixed. There is a unique bounded solution; all other solutions diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
Since $\Re(w)>0$, we must have
$\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|=0,$
which gives $z=\mu(w)$. Now, prove that that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$. By definition of $\mu(w)$, we have
$\mu(w)+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du=-(1-w)\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$
Using equation (2), we then obtain
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
By Hypothesis we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|\leq\frac{|1-w|}{\Re(w)}.$
The differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We start to exhibit this order between solutions through the following proposition.
Proposition 4.
Consider the continuous solutions of the differential equation (1) as given in Notation 1. For every $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ with $\sigma\in(0,1)$, the function $\delta_{s}$ defined as
$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
satisfies the following integral equation:
$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
Denote by $L$ the critical ligne difened as
$L:=\{w\in\mathbb{C},\quad\Re(w)=\frac{1}{2}\}.$
Given the previous proposition, we consider the following rigidity hypothesis:
$\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$ (H)
The previous hypothesis represents a rigidity of the sign along the critical line. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
The following theorem shows that, under the rigidity hypothesis (H), if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 5.
Consider the function $\mu$ given by Notation 2. Under hypothesis (H), for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
$\big(\mu(s),\,\mu(1-\overline{s})\big)\neq(1,1).$
2 Proof of the Proposition 4 and Theorem 5
Proof of the Proposition 4.
Let be $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\Re\big(\psi_{w}(1,t)\big).$ (4)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+t^{-\frac{3}{2}}\Re\big((1-w)\,t^{-i\,\beta}\,\eta(t)\big),\quad x_{w}(1)=1.$ (5)
Let be $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ such that $\sigma\in(0,1)$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\bar{s}}(t)=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\bar{s}}(t)\big)+t^{-\frac{3}{2}}\Re\Big((1-2i\tau)t^{-i\,\tau}\,\eta(t)\Big).$
Use the fact that $x_{s}(1)=x_{1-\bar{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2+\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big).$
From the differential equation (5), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\bar{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}}\Re\big(\,t^{-i\,\tau}\,\eta(t)\big).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2(\sigma-\frac{1}{2})t^{-1}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\,\tau}\,\eta(t)\big)\Big)$
$+(\sigma-\frac{1}{2})t^{-1}\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\,du\Big).$
Use the fact that ${x}_{s}(1)-{x}_{1-\bar{s}}(1)=0$ and integrate, we obtain
${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+2(\sigma-\frac{1}{2})\Re\big(\,\int_{1}^{t}u^{-1}\Big(1-u^{-\frac{1}{2}}u^{-i\,\tau}\,\eta(u)\big)du\Big)$
$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\,du\Big).$
From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
$\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],\quad\forall\beta\in\mathbb{R}.$
Therefore,
${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{3}{2}}\Re\Big(\psi_{\frac{1}{2}+i\,\tau}(1,u)-u^{-i\,\tau}\,\eta(t)\Big)\,du,$
and that can be written as
${x}_{s}(t)-{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big).$
By the notation (4), we have
${x_{s}}(t)-{x}_{1-\bar{s}}(t)=t^{-\frac{1}{2}}\delta_{s}(t):=t^{-\frac{1}{2}}\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big).$
Proof of the Theorem 5.
Let be $s:=\sigma+i\,\tau\in\mathbb{C}$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
$\delta_{s}(t):=\Re\big(\psi_{s}(1,t)-\psi_{1-\bar{s}}(1,t)\big).$
From the equation (1) , we have
$\dot{\delta}_{s}(t)=t^{-1}\Re\Big(\sigma\psi_{s}(1,t)-(1-\sigma)\psi_{1-\bar{s}}(1,t)\Big)-(2\sigma-1)\,t^{-1}\Re\big(t^{-i\,\tau}\eta(t)\Big),$
Since $\psi_{s}(1,1)=\psi_{1-\bar{s}}(1,1)=1$ then ${\delta}_{s}(1)=0$. By hypothesis, we have $|\eta(t)|<1$ for all $t\geq 1$. Since $\sigma\in(\frac{1}{2},1)$ the previous differential equation implies $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
$\delta_{s}(t)>0,\quad\forall t\in(1,1+\epsilon).$
Define the time $t_{*}>1$ as
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ \delta_{s}(u)>0\}.$
To prove of the present Proposition it is equivalent to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$\delta_{s}(t_{*})=0,$ (6)
and we have
$\delta_{s}(u)>0,\quad\forall u\in(1,t_{*}).$ (7)
Thank to the Proposition 4, we have
$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
Since $2\sigma-1>0$, thanks to the hypothesis (H), we obtain
${\delta}_{s}(t)\geq(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t\geq 1.$ (8)
We obtain
${\delta}_{s}(t_{*})\geq(\sigma-\frac{1}{2})^{2}t_{*}^{\frac{1}{2}}\int_{1}^{t_{*}}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du.$
Since $2\sigma-1>0$, by equation (7) we get
${\delta}_{s}(t_{*})>0,$
We obtained a contradiction with equation (6). Then $t_{*}=+\infty$. In other words, $\delta_{s}(t)>0$ for all $t>1$. The equation (8) implies
${\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{2}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t>1.$
Then $\lim_{t\to+\infty}\delta_{s}(t)=+\infty$. This implies that the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ are not both bounded on $1,+\infty)$. By Lemma 3, we deduce that
$(1,1)\neq\big(\mu(s),\,\mu(1-\overline{s})\big).$
Remark 6.
  • Under hypothesis (H), the process used in the proofs of Proposition 4 and Theorem 5 shows that: For all $s\in\mathbb{C}$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following implication
    $\Re\Big(\mu(s)+\mu(1-\overline{s})\Big)\geq 2\implies\Re\Big(\mu(s)-\mu(1-\overline{s})\Big)<0.$
    In this case, the function ${\delta}_{s}$ start with a negative values at 11. Hence, the proof of Theorem 5 cannot be continued.
  • We recall that the rigidity hypothesis (H) is given as:
    $\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$
    From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
    $\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],$
    Hence, the rigidity hypothesis (H) is equivalent to:
    $\forall\beta\in\mathbb{R},\quad\forall t\geq 1:\quad\Re\big(f_{\beta}(t)\big)\geq 0,$
    where
    $f_{\beta}(t):=$ $\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du$
    $+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\eta(v)\,dv\,du.$
3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will see in the following proposition that the hypothesis (H) is also satisfied
Proposition 7.
When $\eta(t)=\{t\}$ the rigidity hypothesis (H) is satisfied.
Proof.
Given the second item of Remark 6, we prove that : For all $\beta\in\mathbb{R}$ and $t\geq 1$ we have $\Re\big(f_{\beta}(t)\big)\geq 0$, where
$f_{\beta}(t):=\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du.$
We have
$f_{\beta}(t)=-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big]\,du,$
equivalent to
$f_{\beta}(t)=\int_{1}^{t}u^{-\frac{1}{2}}h_{\beta}(u)\,du=t^{-\frac{1}{2}}\int_{1}^{t}h_{\beta}(u)\,du+\frac{1}{2}\int_{1}^{t}u^{-\frac{3}{2}}\int_{1}^{u}h_{\beta}(v)\,dv\,du,$
where
$h_{\beta}(u):=-u^{-1-i\,\beta}\,\{u\}+u^{-\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big].$
Then it is sufficient to prove that $\int_{1}^{t}\Re\big(h_{\beta}(u)\big)\,du\geq 0$ for all $t\geq 1$. We have
$\int_{1}^{t}h_{\beta}(u)\,du$ $=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du$
$=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+2(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du-2(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du,$
We obtained
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du$
$-(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du.$
Since $\{u\}=u-\lfloor u\rfloor$, and since
$(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{1}{2}-i\,\beta}\,du-(1-i\beta)\int_{1}^{t}u^{-i\,\beta}\,\,du=0,$
then
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\lfloor u\rfloor\,du$
$+(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\lfloor u\rfloor\,du,$
which can be written as
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1$
$-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\,du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-\frac{3}{2}-i\,\beta}\,du\Big]$
$+(1-i\beta)\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}u^{-1-i\,\beta}\ du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-1-i\,\beta}\,du\Big].$
Integrate, to obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-t^{\frac{1}{2}}\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}\lfloor t\rfloor\big(t^{-\frac{1}{2}-i\,\beta}-\lfloor t\rfloor^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\lfloor t\rfloor\big(t^{-i\,\beta}-\lfloor t\rfloor^{-i\,\beta}\big)$
$-\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-\frac{1}{2}-i\,\beta}-n^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-i\,\beta}-n^{-i\,\beta}\big).$
Using telescopic series, one obtains
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-\frac{1}{2}-i\,\beta}\Big)$
$+\frac{1-i\beta}{i\,\beta}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-i\,\beta}\Big),$
which is equivalent to,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1-i\beta}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
or,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)t^{1-i\,\beta}$
$-\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
Use the Abel summation formula to the last line, we obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}+t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}\,du$
$-\frac{1}{1-i\,\beta}-\int_{1}^{t}u^{-1-i\beta}\{u\}\,du.$
A direct estimate show that $\Re\big(f_{\beta}(t)\big)>0$, for all $t>1$. ∎
Then the rigidity hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\{u\}\,du,\quad\mu(1-\bar{s})=-\bar{s}\int_{1}^{+\infty}u^{-2+\bar{s}}\{u\}\,du.$
In other words, by considering the integral representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have the relations
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
Since the hypothesis (H) is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line. The results provides an answer to the Dynamical Conjecture formulated in [2].